OpenStudy (anonymous):

(u^2+7u/u^2+3u-28)/(u/u+4)

OpenStudy (anonymous):

divide and simplify

OpenStudy (anonymous):

2^(2x) - 3(2^x) = 28 note that (2^x)^2 = 2^(2x) if we let u = 2^x, this becomes u^2 - 3u = 28 u^2 - 3u - 28 = 0 (u - 7)(u + 4) = 0 so either u - 7 = 0 ==> u = 7 or u + 4 = 0 ==> u = -4 but since u = 2^x, then u = -4 is extraneous so u = 7 = 2^x and converting to the log form, x = log_2 7 (or log base 2 of 7)

OpenStudy (anonymous):

I think you confused me

OpenStudy (anonymous):

no never

OpenStudy (anonymous):

isn't is suppose to be x^2 not 2^x

OpenStudy (anonymous):

or is that the same thing

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is same man

OpenStudy (anonymous):

so what would be the fraction form of the answer

OpenStudy (anonymous):

u/u is 1...i dont get it

OpenStudy (anonymous):

Bryan, kindly check that you wrote the question correctly, because as I5starjordan says, u/u = 1. Thanks =)

OpenStudy (anonymous):

use parentheses better